OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 3x - 4y = -12.

OpenStudy (anonymous):

@eric_d

OpenStudy (eric_d):

rearrange them y=mx+b b=y-intercept then substitute value of b to find x-intercept

OpenStudy (anonymous):

uhhh ,,,, @eric_d

OpenStudy (anonymous):

im confused): @eric_d

OpenStudy (anonymous):

@Orion1213

OpenStudy (aum):

3x - 4y = -12 To find the x-intercept, set y = 0 and solve for x. To find the y-intercept, set x = 0 and solve for y.

OpenStudy (anonymous):

the intercept form for a line:\[\frac{x}{a}+\frac{y}{b}=1\]where "a" is the x-intercept and "b" is the y-intercept....